Suppose the utility function for a firm manager is Z=2H1H2 and his income is Y=P1H1 + P2H2. derive the demand equation for H1 and H2.


1
Expert's answer
2021-06-02T12:25:04-0400

Given

Utility function "Z=2H_1H_2"

and budget constraint as "Y= P_1H_1+P_2H_2"

"MRS( H_1H_2)=-\frac{Mu_1}{Mu_2}"

"Mu_1=-H_1\/\/Mu_2=-H_2"

"-\frac{H_1}{H_2}=-\frac{P_1}{P_2}"

Demand for H1

"H_1=\frac{Y}{2P_1}"

Demand for H2

"H_2=\frac{Y}{2P_2}"